NDOS Commissioning Meeting 2011/05/12

Attendance: Jon, Leon, Susan, Zeb, Nick, Peter, Mat, Kurt, Jim, Denis, Minerba, Gary, Chad, Athans


  1. NDOS Status and Channel Recovery Plans
  2. DAQ/Electronics Update & Plans
  3. DCS Update & Plans
  4. DataCheck Update
  5. AOB (All)

NDOS Status

  • Jon:
    • uptime (run>30 minutes long and all 9 dcms in the readout) during the last week was ~66%
    • active channel fraction (whole detector) has been ~77% this past week.
  • Mat:
    • Karen managed to fill ~all obstructed modules. Everything is now topped off and filled except for the module that had leaked recently. That has been patched, but not yet re-filled.
    • No channel recover this week, waiting on parts.
    • Denis and Mat have been trying to catalog APD cooling results.
    • Humidity is high again, Mat will run out to NDSB to check on humidifier and will follow up with Bill Moorehouse if necessary
    • Chad sent an email out about flowmeters... did Leon get the email? Yes.

DAQ/Electronics Update & Plans

  • Peter: message facility moved forward which required new compiler and recompile of all packages.
  • Andrew: still having problems getting TDU in MINOS building to run spillserver. The switches they gave us is incompatible with our DAQ (frame size). Cat 6 copper needs to be run between the MINOS SB and NDSB. Ticket for this was submitted yesterday.

DCS Update & Plans

  • Athans:
    • software development progressing with both AD and CD experts.
    • testing of DCS readout while taking data has been successful
  • Andrew: had a discussion with Jim Patrick after yesterday's DCS meeting. He feels that EPICS is the way to go for the short-term solution. Denis Niklaus and some others could potentially help out with the EPICS.

DataCheck Update

  • Nick presented
  • correlations were search for between data and DCS for a run (12002) that showed changes in data size and rate in the past. No real strong correlations were found for this run.
  • cut studies were done

AOB (All)